# Analysis of Decentralized Stochastic Successive Convex Approximation for composite non-convex problems

### v1

@@ Abstract
Successive Convex approximation (SCA) methods have shown to improve the empirical convergence of non-convex optimization problems over proximal gradient-based methods. SCA uses a strongly convex surrogate and offers a more flexible framework to solve such optimization problems. Further, in decentralized optimization, which aims to optimize a global function using only local information, the SCA framework has been successfully applied to achieve improved convergence. Still, the stochastic first order (SFO) complexity decentralized SCA algorithms have remained under-studied. While non-asymptotic convergence analysis has been studied for decentralized deterministic settings, its stochastic counterpart has only been shown to converge asymptotically.
We have analyzed a novel accelerated variant of the decentralized stochastic SCA that minimizes the sum of non-convex (possibly smooth) and convex (possibly non-smooth) cost functions. The algorithm viz. Decentralized Momentum-based Stochastic SCA (D-MSSCA), iteratively solves a series of strongly convex subproblems at each node using one sample at each iteration. The key step in non-asymptotic analysis involves proving that the average output state vector moves in the descent direction of the global function. This descent allows us to obtain a bound on average iterate progress and mean-squared stationary gap. The recursive momentum-based updates at each node contribute to achieving stochastic first order (SFO) complexity of $\mathcal{O}(\epsilon^{-3/2})$ provided that the step sizes are smaller than the given upper bounds. Even with one sample used at each iteration and a non-adaptive step size, the rate is at par with the SFO complexity of decentralized state-of-the-art gradient-based algorithms. The rate also matches the lower bound for the centralized, unconstrained optimization problems. Through a synthetic example, the applicability of D-MSSCA is demonstrated.
@@ I Introduction
where $u_{i}({\mathbf{x}})={\mathbb{E}}\left[f_{i}({\mathbf{x}},{\boldsymbol{\xi}}_{i})\right]$. Local objective function $f_{i}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is assumed to be non-convex, smooth and known only to agent $i$. Regulariser $h:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is a convex and possibly non-smooth function and constraint $g:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is a convex function. Also, both $g,h$ are publicly known. The form of problem ($\mathcal{P}$) arises in multiple areas, such as statistical inference, decision-making in sensor networks, and machine learning problems [1].
Existing methods to solve ($\mathcal{P}$) include projected and proximal stochastic gradient methods [2, 3, 4, 5, 6, 7] and Successive Convex Approximation (SCA) [8, 9] methods. The performance of these algorithms is measured in terms of the number of stochastic first order (SFO) oracle calls required to reach a $\epsilon$- Karush-Kuhn Tucher (KKT) point. While [2] proposes a projected DSGD-type algorithm for problems with a compact constraint set, [3] goes ahead and establishes the asymptotic convergence of DSGD for a family of non-convex, non-smooth functions. Further, in [4], a decentralized stochastic proximal primal-dual method called SPPDM is proposed, assuming that the epigraph of $h$ is a polyhedral set. Only three works address non-asymptotic iteration complexity analysis for stochastic non-convex composite problems with a general convex non-differentiable regularizer $h$. While DProxSGT[6] achieves sub-optimal rate of $\mathcal{O}(\epsilon^{-2})$ without mean-squared smoothness assumption, ProxGT-SR-O/E [5] and DEEPSTORM [7] achieve an optimal convergence rate of $\mathcal{O}(\epsilon^{-3/2})$. However, the problem with ProxGT-SR-O/E [5] is that it uses large batches and more communication rounds at each iteration. Even though DEEPSTORM [7] overcomes the use of more communications rounds, it still uses small batches to achieve the optimal rate.
@@
Under decentralized settings, there are only a handful of SCA algorithms [8, 24, 9] including both deterministic and stochastic cases. In [24], the authors proposed an SCA-based decentralized algorithm, NEXT, . A stochastic variant, S-NEXT, was proposed in [9]; however, they did not apply the momentum to the update steps. In both works, only asymptotic convergence has been proven. Recently, [8] introduced a decentralized momentum-based algorithm that employs Nesterov-like momentum, providing the first non-asymptotic analysis of decentralized SCA methods for deterministic case. However, their analysis introduced a new metric tailored to SCA, and the detailed proofs were only provided for the case where the functions $u_{i}$-s are convex. To the best of our knowledge, there is no comprehensive non-asymptotic convergence analysis for decentralized consensus stochastic non-convex problems within the SCA rubric in the literature.
In this work, we have analyzed a novel Decentralized Momentum-based Stochastic SCA (D-MSSCA) algorithm to solve decentralized stochastic non-convex composite optimization problems. The D-MSSCA hinges on SCA techniques and iteratively solves a convexified subproblem at each node using recursive momentum [25] type local gradient updates to reach the $\epsilon-$KKT point with an optimal convergence rate of $\mathcal{O}(\epsilon^{-3/2})$. Our analysis extends the methods used in gradient-based approaches [5, 7] to SCA framework. The key challenge in the convergence analysis was to form a descent inequality for the global function $U$ at the average state-vector of D-MSSCA in terms of iterate progress, which we have overcome by utilizing the strong convexity of the surrogate. This descent helped us in bounding the mean-squared stationary gap, a more general metric than that used in [8] for analyzing non-convex functions. Unlike the Nesterov updates used in [8], our use of recursive momentum-type updates significantly contributes to this advancement. Finally, simulations on a synthetic problem empirically validate the theoretical findings.
A comparative performance of various state-of-the-art algorithms that can be used to solve ($\mathcal{P}$) has been provided in Table I. It can be observed that even with one sample per iteration, the proposed D-MSSCA algorithm is able to achieve the optimal convergence rate. Additional remarks are also provided in the table.
@@ I-A Notations
We denote vectors (matrices) using lowercase (uppercase) bold font letters. For a vector ${\mathbf{x}}$, we denote its transpose by ${\mathbf{x}}^{\mathsf{T}}$ and its $i$-th element by $[{\mathbf{x}}]_{i}$. Likewise, the $(i,j)$-th component of ${\mathbf{A}}$ is given by $A_{ij}$. An $n$-dimensional identity matrix is denoted by ${\mathbf{I}}_{n}$. The $n$-dimensional all-one vector is denoted by $\mathbf{1}_{n}$. The Kronecker product is denoted using $\otimes$. The $d$-dimensional average of any $nd-$dimensional vector $\mathbf{a}\in{\mathbb{R}}^{nd}$, is represented by $\bar{\mathbf{a}}=\frac{1}{n}\left(\mathbf{1}_{n}^{\mathsf{T}}\otimes{\mathbf{I}}_{d}\right)\mathbf{a}\in{\mathbb{R}}^{d}$. For a vector ${\mathbf{y}}$, $\left\|{\mathbf{y}}\right\|$ denotes its $\ell_{2}$ (Euclidean) norm, Maximum eigenvalue of ${\mathbf{A}}$ is represented by $\lambda_{\max}({\mathbf{A}})$. The set of neighbours of node $i$ is denoted by ${\mathcal{N}}_{i}$. For a non-smooth function $h+11_{{\mathcal{X}}}$, $\partial\left(h+11_{{\mathcal{X}}}\right)\mid_{{\mathbf{x}}=\mathbf{a}}$ represents the set of sub-gradients of $h(x)+11_{{\mathcal{X}}}(x)$ at ${\mathbf{x}}=\mathbf{a}$. For the ease of writing we have defined $\nabla_{x}{\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})\mid_{{\mathbf{x}}=\mathbf{a}}:={\boldsymbol{\nabla}}{\tilde{f}}(\mathbf{a},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})$.
The rest of the paper is organized as follows. Section II discusses the proposed algorithm, and the assumptions on the problem are discussed. In Section III, the convergence proof of the proposed algorithm is presented. Section IV briefly describes the proposed algorithm’s applicability to a synthetic problem.
@@
ProxGT-SR-O/E [5] $\mathcal{O}(\epsilon^{-3/2})$ Non-Asymptotic Multiple communication, larger batches per iteration
@@
DEEPSTORM [7] $\mathcal{O}(\epsilon^{-3/2})$ Non-Asymptotic gradient-based, small batches per iteration
@@ II-B Proposed Algorithm
We will now state the proposed algorithm. Each node is initialized at an arbitrary feasible point ${\mathbf{x}}_{i}^{1}\in{\mathcal{X}}$, which satisfies $g({\mathbf{x}}_{i}^{1})\leq 0$. Each node $i$ constructs a strong convex surrogate ${\hat{f}}_{i}$ and solves the following optimization problem based upon the private knowledge of $f_{i}$ and public knowledge of $h$ and $g$;
${\hat{{\mathbf{x}}}}_{i}^{t}=\underset{{\mathbf{x}}_{i}\in\mathcal{X}}{\arg\min}$ $~{}~{}{\tilde{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)+\pi_{i}^{t}\left({\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\right)+h({\mathbf{x}}_{i})$ (1)
@@
${\tilde{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)={\hat{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)+(1-\beta)\left({\mathbf{z}}_{i}^{t-1}-\nabla f_{i}({\mathbf{x}}_{i}^{t-1},\xi_{i}^{t})\right)\left({\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\right)$ (2)
with $\pi_{i}^{t}={\mathbf{y}}_{i}^{t}-{\mathbf{z}}_{i}^{t}$ and $\beta$ is the step size. One choice of ${\hat{f}}$ can be ${\hat{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)=f_{i}\left({\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)+\nabla f_{i}\left({\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)\left({\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\right)+\frac{\mu}{2}\left\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\right\|^{2}$. Each node then performs the following two updates
@@
This update is inspired from [26]. It should be noted that in (4) we have used local momentum-based gradient estimator ${\mathbf{z}}_{i}^{t}$ [25, 27]. Also, it is noteworthy that the update (4) can also be seen as a convex combination of vanilla SGD and SARAH-type gradient estimator [28]. Finally, we perform a global gradient update:
@@ II-C1 Assumptions on ($\mathcal{P}$)
$U$ is bounded below, i.e., $\underset{{\mathbf{x}}\in{\mathbb{R}}^{d}}{\inf}U(x)>-\infty$
@@
Each local function $f_{i}$ is $L_{i}-$ smooth;
${\mathbb{E}}\left\|\nabla f_{i}({\mathbf{x}}^{t},\xi^{t})-\nabla f_{i}({\mathbf{y}}^{t},\xi^{t})\right\|=L_{i}{\mathbb{E}}\left\|{\mathbf{x}}^{t}-{\mathbf{y}}^{t}\right\|;$
Global function $u$ is $L-$smooth, and $L_{\max}=\max_{i}\{L_{i}\}$, where $\sum_{i=1}^{n}L_{i}\leq L\leq nL_{\max}$
Assumptions A1- A4 are standard in the context of distributed optimization [29]. A direct consequence of A1 is that for an initial point ${\mathbf{x}}_{i}^{1}\in{\mathcal{X}}$ we have $U({\bar{{\mathbf{x}}}}^{1})-U^{\star}\leq B_{1}$. Assumption A4 implies that $U$ is also $L$-smooth. is introduced to simplify the analysis.
@@ II-C2 Assumptions on the surrogate
Each surrogate ${\hat{f}}_{i}$ of local function $f_{i}$ is $\mu_{i}-$ strongly convex.

### v2

@@ Abstract
Successive Convex approximation (SCA) methods have shown to improve the empirical convergence of non-convex optimization problems over proximal gradient-based methods. SCA uses a strongly convex surrogate and offers a more flexible framework to solve such optimization problems. Further, in decentralized optimization, which aims to optimize a global function using only local information, the SCA framework has been successfully applied to achieve improved convergence. Still, the stochastic first-order (SFO) complexity of decentralized SCA algorithms has remained understudied. While non-asymptotic convergence analysis has been studied for decentralized deterministic settings, its stochastic counterpart has only been shown to converge asymptotically.
We have analyzed a novel accelerated variant of the decentralized stochastic SCA that minimizes the sum of non-convex (possibly smooth) and convex (possibly non-smooth) cost functions. The algorithm viz. Decentralized Momentum-based Stochastic SCA (D-MSSCA), iteratively solves a series of strongly convex subproblems at each node using one sample at each iteration. The recursive momentum-based updates at each node contribute to achieving stochastic first order (SFO) complexity of $\mathcal{O}(\epsilon^{-3/2})$ provided that the step sizes are smaller than the given upper bounds. Even with one sample used at each iteration and a non-adaptive step size, the rate is at par with the SFO complexity of decentralized state-of-the-art gradient-based algorithms. The rate also matches the lower bound for the centralized, unconstrained optimization problems. Through a synthetic example, the applicability of D-MSSCA is demonstrated.
@@ I Introduction
where $u_{i}({\mathbf{x}}):=\mathbb{E}[f_{i}({\mathbf{x}},{\boldsymbol{\xi}}_{i})]$ with $f_{i}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ being smooth but possibly non-convex. On the other hand, $h:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is convex but possibly non-smooth while $g:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is a convex function. The agents communicate over undirected links, denoted by tuples of the form $(i,j)$. The topology of the communication network is represented by an undirected graph $\mathcal{G}=({\mathcal{V}},{\mathcal{E}})$, where the node set ${\mathcal{V}}$ collects the set of agents and the edge set ${\mathcal{E}}$ collects the set of communication links.
We consider the fully decentralized setting, where $f_{i}$ is private to agent $i$, though the regularizer $h$ and the constraint function $g$ are public knowledge. Nevertheless, the agents are unaware of the global objective function but must communicate with each other in order to solve ($\mathcal{P}$). The decentralized stochastic optimization problem in ($\mathcal{P}$) arises in a number of areas, including statistical inference, machine learning, and sensor networks [1]. In all of these applications, the objective $f_{i}$ depends on the data or observation ${\boldsymbol{\xi}}_{i}$ which is private to agent $i$ and cannot be shared with the other agents. Instead, the agents may only share the derived quantities such as gradients or iterates.
Existing methods to solve ($\mathcal{P}$) include projected and proximal stochastic gradient methods [2, 3, 4, 5, 6, 7] and Successive Convex Approximation (SCA) [8, 9] methods. The performance of these algorithms is measured in terms of the number of stochastic first order (SFO) oracle calls required to reach a $\epsilon$-Karush-Kuhn Tucher (KKT) point. While [2] proposes a projected DSGD-type algorithm for problems with a compact constraint set, [3] goes ahead and establishes the asymptotic convergence of DSGD for a family of non-convex, non-smooth functions. Further, in [4], a decentralized stochastic proximal primal-dual method called SPPDM is proposed, assuming that the epigraph of $h$ is a polyhedral set. Only three works address non-asymptotic iteration complexity analysis for stochastic non-convex composite problems with a general convex non-differentiable regularizer $h$. While DProxSGT[6] achieves sub-optimal rate of $\mathcal{O}(\epsilon^{-2})$ without mean-squared smoothness assumption, ProxGT-SR-O/E [5] and DEEPSTORM [7] achieve an optimal convergence rate of $\mathcal{O}(\epsilon^{-3/2})$. However, the problem with ProxGT-SR-O/E [5] is that it uses large batches.
@@
Under decentralized settings, there are only a handful of SCA algorithms [8, 24, 9] including both deterministic and stochastic cases. Authors in [24, 9], proposed an SCA-based decentralized algorithm NEXT, and its stochastic variant S-NEXT, [9]. In both of these works, only asymptotic convergence has been proven. Recently, [8] introduced a decentralized algorithm that employs Nesterov-like momentum, providing the first non-asymptotic analysis of decentralized SCA methods for deterministic case. However, the rate metric used is not general but SCA-specific, and detailed proofs were only provided for the case where the objective function is convex. To the best of our knowledge, there is no comprehensive non-asymptotic convergence analysis for the decentralized stochastic SCA algorithm in the literature.
In this work, we have analyzed a novel Decentralized Momentum-based Stochastic SCA (D-MSSCA) algorithm to solve ($\mathcal{P}$). The D-MSSCA hinges on SCA technique and iteratively solves a convexified subproblem at each node. The recursive momentum type local gradient estimate along with global gradient tracking allowed us to achieve an optimal convergence rate of $\mathcal{O}(\epsilon^{-3/2})$. Our analysis extends the methods used in gradient-based approaches [5, 7] to SCA framework. One of the challenges in the convergence analysis is obtaining the global descent direction using local or partial information. We overcome this by using the local optimality condition of the strongly convex subproblem at each node. This, in turn, allowed us to develop a recursive bound on the global function $U$ characterizing the progress in a single iteration. Finally, simulations on a synthetic problem empirically validate the theoretical findings.
A comparative performance of various state-of-the-art algorithms that can be used to solve ($\mathcal{P}$) is provided in Table I. It can be observed that the proposed D-MSSCA algorithm achieves the optimal convergence rate.
@@ I-A Notations
We denote vectors (matrices) using lowercase (uppercase) bold font letters. The $i$-th entry of vector ${\mathbf{x}}$ is denoted by $[{\mathbf{x}}]_{i}$ while the $(i,j)$-th entry of matrix ${\mathbf{A}}$ is denoted by $A_{ij}$. The $n\times n$ identity matrix is denoted by ${\mathbf{I}}_{n}$ while the $n\times 1$ all-one vector is denoted by $\mathbf{1}_{n}$. The Kronecker product is denoted by $\otimes$. The $d$-dimensional average of any $nd$-dimensional vector $\mathbf{a}\in{\mathbb{R}}^{nd}$, is represented by $\bar{\mathbf{a}}=\frac{1}{n}\left(\mathbf{1}_{n}^{\mathsf{T}}\otimes{\mathbf{I}}_{d}\right)\mathbf{a}\in{\mathbb{R}}^{d}$. The Euclidean norm of ${\mathbf{y}}$ is denoted by $\left\|{\mathbf{y}}\right\|$. The maximum eigenvalue of a matrix ${\mathbf{A}}$ is denoted by $\lambda_{\max}({\mathbf{A}})$. The set of subgradients of a function $v:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ calculated at ${\mathbf{x}}={\mathbf{a}}$ is denoted by $\partial\left(v\right)\mid_{{\mathbf{x}}={\mathbf{a}}}$. For the sake of brevity, we define $\nabla_{\mathbf{x}}{\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})\mid_{{\mathbf{x}}=\mathbf{a}}:={\boldsymbol{\nabla}}{\tilde{f}}(\mathbf{a},{\mathbf{x}}_{i}^{t},{\boldsymbol{\xi}}_{i}^{t})$. The indicator function is defined as $11_{{\mathcal{X}}}({\mathbf{x}})$ where $011_{{\mathcal{X}}}({\mathbf{x}})=$ 0 if ${\mathbf{x}}\in{\mathcal{X}}$, otherwise $11_{{\mathcal{X}}}({\mathbf{x}})=\infty$. Finally, ${\mathcal{H}}^{t}$ represents the history of the system generated by $\{\xi_{i}^{\tau}\}_{i=\{1,2,..,n\}}^{\tau\leq t-1}$.
The rest of the paper is organized as follows: Section II discusses the proposed algorithm and the assumptions on the problem are discussed. In Section III, the convergence proof of the proposed algorithm is presented. Section IV briefly describes the proposed algorithm’s applicability to a synthetic problem.
@@
ProxGT-SR-O/E [5] $\mathcal{O}(\epsilon^{-3/2})$ Non-Asymptotic Multiple communication, large batches per iteration
@@
DEEPSTORM [7] $\mathcal{O}(\epsilon^{-3/2})$ Non-Asymptotic gradient-based
@@ II-B Proposed Algorithm
We will now state the proposed algorithm. All nodes are initialized at the same feasible point ${\bar{{\mathbf{x}}}}^{1}\in{\mathcal{X}}$, which satisfies $g({\bar{{\mathbf{x}}}}^{1})\leq 0$. Each node $i$ constructs a strong convex surrogate ${\hat{f}}_{i}$ and solves the following optimization problem based upon the private knowledge of $f_{i}$ and public knowledge of $h$ and $g$;
${\hat{{\mathbf{x}}}}_{i}^{t}=\underset{{\mathbf{x}}_{i}\in\mathcal{X}}{\arg\min}$ $~{}~{}{\tilde{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)+\langle\pi_{i}^{t},{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\rangle+h({\mathbf{x}}_{i})$ (1)
@@
${\tilde{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)={\hat{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)+(1-\beta)\langle{\mathbf{z}}_{i}^{t-1}-\nabla f_{i}({\mathbf{x}}_{i}^{t-1},\xi_{i}^{t}),{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\rangle$ (2)
with $\pi_{i}^{t}={\mathbf{y}}_{i}^{t}-{\mathbf{z}}_{i}^{t}$ and $\beta$ is the step size. One choice of ${\hat{f}}$ can be ${\hat{f}}_{i}\left({\mathbf{x}}_{i},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)=f_{i}\left({\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)+\nabla f_{i}\left({\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)\left({\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\right)+\frac{\mu}{2}\left\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\right\|^{2}$. Each node then performs the following two updates,
@@
This update is inspired from [25]. It should be noted that in (4) we have used local momentum-based gradient estimator ${\mathbf{z}}_{i}^{t}$ [26, 27]. Also, it is noteworthy that the update (4) can also be seen as a convex combination of vanilla SGD and SARAH-type gradient estimator [28]. Finally, using gradient tracking[29, 30], each node updates the local estimate of the global gradient:
@@ II-C1 Assumptions on ($\mathcal{P}$)
The objective $U$ is bounded below, i.e., $\underset{{\mathbf{x}}\in{\mathbb{R}}^{d}}{\inf}U(x)>-\infty$
@@
Mean Squared Smoothness(MSS): Each local function $f_{i}$ is $L$-smooth ,
${\mathbb{E}}\left\|\nabla f_{i}({\mathbf{x}}^{t},\xi^{t})-\nabla f_{i}({\mathbf{y}}^{t},\xi^{t})\right\|=L{\mathbb{E}}\left\|{\mathbf{x}}^{t}-{\mathbf{y}}^{t}\right\|;$
Assumptions A1- A4 are standard in the context of distributed optimization. A direct consequence of A1 is that for an initial point ${\mathbf{x}}_{i}^{1}\in{\mathcal{X}}$ we have $U({\bar{{\mathbf{x}}}}^{1})-U^{\star}\leq B$. Assumption A4 implies that $u$ is also $L$-smooth. is introduced to simplify the analysis.
@@ II-C2 Assumptions on the surrogate
Each surrogate ${\hat{f}}_{i}$ of local function $f_{i}$ is $\mu-$ strongly convex.
@@ II-C3 Assumption on Network
Assumption A7 is standard in the context of decentralized optimization and is required to achieve consensus among agents.